OpenStudy (anonymous):

if h(x)=(fºg)(x) and h(x)=sqrtx+5, find g(x) of f(x)=sqrt x+2

OpenStudy (shamil98):

you mean √x?

OpenStudy (anonymous):

yes square root

OpenStudy (shamil98):

okay, (f o g) means you plug the value of g into f.

OpenStudy (anonymous):

ok?

OpenStudy (anonymous):

g= x + 9

OpenStudy (anonymous):

really?

OpenStudy (anonymous):

ya

OpenStudy (anonymous):

so g(x)=x+9

OpenStudy (anonymous):

ya

OpenStudy (anonymous):

x+3 was right =(

OpenStudy (anonymous):

no

OpenStudy (***[isuru]***):

@880447017 explain how ya got ( x+9) ?

OpenStudy (***[isuru]***):

and... @brucek69 ur values... they r not clear \[h(x) = \sqrt{x + 5} \] \[f(x) = \sqrt{x + 2}\] ????

OpenStudy (***[isuru]***):

r these the expressions ya got ? am I correct ?